OpenStudy (anonymous):

Please check this question. Jimmy is comparing two jobs he was offered. Company 1 has offered him a payment package of $400 a month with an 8 percent commission rate on any sale he makes. Company 2 offered a graduated package commission of 5 percent on the first $4000 of sales and 9 percent commission on any sales above that. If he sells $12000 worth of merchandise in a month, which company will pay him the most money?

OpenStudy (anonymous):

@hba

OpenStudy (anonymous):

Be right back. 5 minutes.

OpenStudy (anonymous):

Let M = Months, C = commission. For one month Company A: 400*M+0.08*12000 = 400(1) + 960 = 1360 Company B: 0.05*4000 + 0.09*8000 = 200+720 = 920 Therefore Company A

OpenStudy (anonymous):

I got that as well. Thanks.

OpenStudy (anonymous):

I wanted to check wether or not i was right.

OpenStudy (anonymous):

No problem...you're most welcome

OpenStudy (anonymous):

:)